The president of Hill Enterprises, Terri Hill, projects the firms aggregate demand requirements over the next 8 months as foa) the total cost of hiring=$ (enter your response as a whole number)

b)the total cost of layoffs=$?(enter your response as a whole number)

c)the total inventory carrying cost =$?(enter your response as a whole number)

d)the total stockout cost$?(enter your response as a whole number)

e) the total cost, excluding normal time labor cost, is =$ ?(enter your response as a whole number)
Answer #1

Below is the updated plan and costing based on details given:

Period Month Demand Production Hire (Units) Layoff (units) Ending Inv. Stockouts (units)
0 December             1,600           1,600 - 0 200 -
1 January             1,200           1,600 - 0 200+1600-1200=600 -
2 February             1,700           1,200 - 1600-1200=400 600+1200-1700=100 -
3 March             1,600           1,700 1700-1200=500 - 100+1700-1600=200 -
4 April             1,800           1,600 1700-1600=100 200+1600-1800=0 -
5 May             2,300           1,800 1800-1600=200 - - 2300-1800=500
6 June             2,100           2,300 2300-1800=500 - 2300-2100=200
7 July             1,700           2,100 2300-2100=200 200+2100-1700=600
8 August             1,700           1,700 - 2100-1700=400 600 -
Total 1200 1100 2300-200=2100 500
Cost per unit $55 $75 $20 $125
Total Cost $66,000 $82,500 $42,000 $62,500
$253,000

a) the total cost of hiring=$66,000

b)the total cost of layoffs=$82,500

c)the total inventory carrying cost =$42,000

d)the total stockout cost=$62,500

e) the total cost, excluding normal time labor cost, is =$253,000


Please give thumbs up/ likes if you find this answer helpful. Thank you!

                    ◢